OpenStudy (anonymous):

how to solve √6x+5=√53

OpenStudy (maheshmeghwal9):

As in ur question @nishia if \[\sqrt{a}=\sqrt{b}.\]Then it means that a=b so as in ur question 6x+5=53.

OpenStudy (maheshmeghwal9):

just solve now:D

OpenStudy (anonymous):

im really confused i really need help i just need to get the steps of solving the answer

OpenStudy (maheshmeghwal9):

I have stated above wt the question means then ur question reduces to a linear equation as follows \[\LARGE{6x+5=53.}\] so solve it & get 'x'.

OpenStudy (maheshmeghwal9):

can't u solve linear equations @nishia ????????????

OpenStudy (maheshmeghwal9):

tell me:D I will help u

OpenStudy (anonymous):

x=8

OpenStudy (maheshmeghwal9):

yeah @nishia 100% correct:D very nice!!!!!

OpenStudy (maheshmeghwal9):

Any more doubt now?

OpenStudy (maheshmeghwal9):

tell me if ok:)

OpenStudy (anonymous):

no :)

OpenStudy (maheshmeghwal9):

:)